# Concomitant Entanglement and Control Criticality Driven by Collective Measurements

### v1

@@ Abstract
Adaptive quantum circuits—where a quantum many-body state is controlled using measurements and conditional unitary operations—are a powerful paradigm for state preparation and quantum error correction tasks. They can support two types of nonequilibrium quantum phase transitions: measurement-induced transitions between volume- and area-law-entangled steady states and control-induced transitions where the system falls into an absorbing state or, more generally, an orbit visiting several absorbing states. Within this context, nonlocal conditional operations can alter the critical properties of the two transitions and the topology of the phase diagram. Here, we consider the scenario where the measurements are nonlocal, in order to engineer efficient control onto dynamical trajectories. Motivated by Rydberg-atom arrays, we consider a locally constrained model with global sublattice magnetization measurements to steer the system’s dynamics onto a many-body orbit with finite recurrence time. With the aid of a suitable classical limit, we diagnose the control transition to be in a nonequilibrium universality class with dynamical exponent $z<1$ that persists upon reintroducing quantum fluctuations. In the quantum limit, an entanglement transition additionally emerges that coincides with the control transition—to within our numerical resolution. Both transitions exhibit a dynamical criticality consistent with recent results on measurement-induced phase transitions in power-law interacting circuits. We attribute this feature and the apparent coincidence of the control and entanglement transitions to the global nature of the control.
@@ I Introduction
With experimentally accessible noisy intermediate scale quantum (NISQ) devices, it is a timely and fundamental question to extend these ideas to monitored quantum systems. In a prototypical class of dynamical maps, Refs [3, 4] defined a probabilistic control protocol for the $\beta$-adic Rényi maps [6] which can also be used to control analogous quantum dynamics [7, 8, 9, 10]. The quantum control problem is enriched by measurements (necessary for the control protocol) which entail a backaction on the quantum state of the system, resulting in a phase transition in the late-time entanglement content of the quantum state. The entanglement transition can either precede or coincide with the (putatively) classical control transition depending on details of the entangling gates and control operation [11, 8, 10]. Thus, probabilistic control of quantum systems fits into the broader context of adaptive quantum dynamics [12, 13, 14, 15, 16, 17, 18, 19, 20, 11, 21, 22, 23, 24, 25, 26], where measurement and feedback can drive critical phenomena separately witnessed by local order parameters and entanglement measures [27, 28, 29, 30, 31, 32, 33]. This class of quantum dynamics holds promise for applications in quantum technologies ranging from robust state preparation to quantum error correction.
@@
The majority of monitored quantum circuits considered to date consist of short- to long-ranged unitary quantum gates and random local measurements. While measurements that project each local qubit fully destroy the entanglement of the many-body wavefunction, it is possible to perform a collective measurement of a global property of the ensemble of spins (e.g., their total magnetization) which has an extensive number of possible measurement outcomes and does not fully disentangle the state. Long-range gates have a strong effect on the universality class of the measurement induced phase transition (MIPT) [34, 35, 36], but it remains unclear if such a transition can persist in the presence of long-range or collective measurements and if so, what is its universal nature. At the same time, adding feedback on top of collective measurements opens the door to a much wider range of possible feedback operations and fixed point structures that can be used to overcome the strongly entangled dynamics.
@@
In this paper, we introduce a model of probabilistic quantum control inspired by Rydberg-atom quantum simulators [37, 38]. The model consists of a chaotic circuit driven by kinetically constrained Rydberg-blockade dynamics [39, 40, 41, 42] that is stochastically interleaved with collective measurements and local control protocol (see Fig. 1). The chaotic circuit (defined by $p=0$) possesses an unstable Hilbert-space orbit where the system cycles through a finite number of product states [43, 44, 39]. The control protocol attempts to convert this unstable orbit into an attractor, i.e., a steady state that is ultimately reached regardless of the initial state, using collective measurements and local correction operations. We demonstrate the existence of control and entanglement transitions in this model with the aid of a classical limit in which the model maps onto a classical probabilistic cellular automaton (CA); this limit manifests a many-body generalization of classical probabilistic control. With a variety of metrics, we estimate the locations and critical exponents of both transitions and find evidence that the transitions coincide and belong to the same universality class (although we cannot rule out the possibility that the transitions occur quite close to one another and with very similar critical exponents). Intriguingly, both transitions possess the same dynamical critical exponent $z\approx 0.88<1$, similar to MIPTs observed in monitored quantum dynamics with long-range unitary gates [34], and share a correlation length exponent $\nu\approx 2$.
@@
The remainder of the paper is organized as follows. In Sec. II.1, we define the probabilistic control model and its classical limit. In Sec. III, we study the control transition in the quantum model and show where its critical properties match the transition in the classical model. In Sec. IV, we study the entanglement transition through the lens of entanglement entropy, mutual information, and purification dynamics. We conclude in Sec. V with a discussion and outlook. Appendices A and B discuss generalizations of our control protocol and the impact of finite-size effects, respectively.
@@ II.2 Control Protocol
Note that this measurement does not collapse |$\ket{\psi}$ to a single CB state. In fact, the subspace $\mathcal{H}_{m_{\rm even(odd)}}$ contains exponentially many CB states and therefore can support volume-law entanglement. For this reason, we refer to this type of measurement, which reveals global information about the state without fully collapsing it to a product state, as a collective measurement. Such a measurement can be implemented experimentally in a qubit-only setup by computing the magnetization onto a register of $O(\log_{2}L)$ ancilla qubits [52, 53], or in a circuit/cavity QED setup by dispersively coupling the qubits in each sublattice to a separate cavity mode and measuring the frequency shift of either mode.
@@
is applied on site $k=j$ or $j+2$. $U^{c}_{k}$ flips qubit $k$ provided the flip does not violate the Fibonacci constraint described below Eq. (1). The table at the bottom of Fig. 3(b) shows whether qubit $j$ or $j+2$ is flipped, depending on the target state and the outcome of the $Z_{j}$ measurement. In Appendix A, we discuss how this control protocol can be generalized to other Hilbert space orbits with finite recurrence times, of the sort discussed in Ref. [39].
The strength of the control is quantified by two parameters: the control rate $p$ that sets the probability with which control is applied at a given time step, and the probability $q$ that sets the fraction of sites to which the local feedback operation is applied. The addition of the “spatial” probability $q$ is particularly important in that it represents a step beyond the probabilistic control literature [3, 2, 4, 5] that has so far focused on single-body chaotic dynamics. In generalizing to the many-body setting, we allow for the possibility that only a finite fraction of sites are subjected to control (beyond the initial step in which some global information about the system’s state is measured).
@@ II.3 Classical Limit
Both the chaotic circuit and the control protocol have a numerically tractable classical limit that is useful in our analysis of the control transition. In this limit, the dynamics of an initial CB state does not lead to superposition and entanglement. Rather, the system is always in a CB state |$\ket{b}$, corresponding to a bitstring $b$. At each time step, the CB state |$\ket{b}\mapsto\ket{b^{\prime}}$ up to a global phase. Thus, to keep track of the dynamics of an initial CB state and the evolution of observables that are diagonal in the CB [like the control order parameter $H_{\rm ZZ}$, Eq. (3)], it is sufficient to follow the dynamics of the bitstring $\dots\mapsto b\mapsto b^{\prime}\mapsto\dots$. The dynamics in the classical limit is therefore equivalent to a CA. This is clearly visible in Fig. 2(b,d,f) where the system always has a fixed expectation value $\braket{Z_{j}}$ on each site, indicating that the system remains in a CB state throughout the dynamics.
@@ III Control Transition
We now consider the existence of a control transition in this model, before turning our focus in Sec. IV to the interplay of control and entanglement properties. Previous work on $\beta$-adic Rényi circuits [7, 8, 9, 10] suggests that properties of the control transition are primarily governed by classical physics. For this reason, we first consider the automaton limit of our model [$\theta=\pi/2$ in Eq. (4)], whose classical simulability gives access to larger system sizes and better estimates of critical properties that can serve as a useful guide when analyzing the quantum limit. In our numerical simulations, we set $q=0.2$ and measure the control order parameter $H_{ZZ}$ after evolving the system for $L$ time steps from an initial CB state. Results are averaged over $10^{4}$ samples comprising randomly chosen initial CB states, circuit realizations, and quantum measurement outcomes (sampled according to their Born probabilities). Each sample yields an expectation value of the control order parameter that we denote $\braket{H_{ZZ}(t)}=\langle\psi(t)|H_{ZZ}|\psi(t)\rangle$. In the classical limit, $\braket{H_{ZZ}(t)}$ takes discrete values between 00 and $1$ in steps of $1/L$ since the system is always in a CB state; in the quantum limit, the system can be in a superposition of CB states and $\braket{H_{ZZ}(t)}$ becomes a continuous variable. The sample average $\overline{\braket{H_{ZZ}(t)}}$, along with error bars denoting the standard error of the mean, is the object of our analysis below.
@@
Before proceeding, we offer a remark about finite-size effects. The control protocol defined in Sec. II.2 targets the vacuum orbit Eq. (2), which is invariant under the chaotic circuit defined in Sec. II.1. If the control protocol successfully drives the system onto the orbit, it can never escape. In a finite-size system, even at low control rate there is a finite probability that the system becomes trapped on the orbit. This probability is exponentially small in the system size $L$, but this means that such rare events can occur on timescales $\sim e^{O(L)}$. At large system sizes like those accessible in the classical limit, this is not an issue. However, at small system sizes like those accessible in the quantum limit, this finite-size effect manifests as a late-time decay of the control order parameter, even in the chaotic phase where one would expect it to saturate to a time-independent value $\overline{\braket{H_{ZZ}(t\rightarrow\infty)}}_{p=0}$. We discuss the decay further in Appendix B and provide a mechanism to remove it if desired. Nevertheless, we will see that it is still possible to perform a consistent finite-size scaling analysis in the presence of this decay  [^1] We note that this effect is also present in studies of absorbing state transitions in monitored quantum dynamics [15, 17, 16], but was not present in studies of probabilistic control in $\beta$-adic Rényi circuits [7, 8, 9, 10], where a random perturbation was used to remove the orbits while preserving the control transition..
@@ III.1 Classical Control Transition
Figure 4: Order parameter (classical model). Shown at time $t=L$ as a function of control rate $p$ in the classical limit of the model for both large (blue) and small (red) system sizes. Here and in all subsequent figures, error bars indicating the standard error of the mean over $10^{4}$ samples are present, but are smaller than the plot markers. Inset: Finite-size scaling collapse of the large-system data indicates $p_{c}=0.49(1)$, $\beta=0.16(2)$, and $\nu=2.1(2)$. Small-system data are collapsed assuming a scaling function of the form $L^{-\beta/\nu}f[L^{1/\nu}(p-p_{c})+AL^{-\alpha}]$ with $p_{c}$, $\beta$, and $\nu$ fixed by the large-system data collapse. This yields $A\approx 9.46$ and $\alpha\approx 0.94$, which quantify the finite-size effects present at system sizes accessible to the quantum numerics.
@@
In Fig. 4, we plot the sample-averaged order parameter $\overline{\braket{H_{ZZ}(t=L)}}$ for a range of $p$ and two families of system sizes: $L=300,\dots,540$ (blue points) and $L=16,\dots,32$ (red points). The large-system-size data will be used to extract the location and critical exponents of the control transition. The small-system-size data represent system sizes accessible to our fully quantum simulations that will be reported in the next subsection. We first describe the transition from the perspective of the large-system data before returning to the small-system data to analyze the finite-size effects that will be important when considering the quantum data.
@@
To extract the critical point $p_{c}$, we perform finite-size scaling collapse assuming a scaling ansatz $\overline{\braket{H_{ZZ}(t/L^{z}\to\infty)}}=L^{-\beta/\nu}f[L^{1/\nu}(p-p_{c})]$, where $z$ is the dynamical critical exponent that will be determined later. This corresponds to a second-order phase transition where the order parameter vanishes at the transition as $\overline{\braket{H_{ZZ}(t/L^{z}\to\infty)}}_{p=p_{c}}\sim|p-p_{c}|^{\beta}$ (in the thermodynamic limit) and the correlation length diverges as $\xi\sim|p-p_{c}|^{-\nu}$. The scaling collapse (like all others reported in this paper) is performed using a standard $\chi^{2}$ analysis [55] that uses a numerical least-squares opimization to collapse the curves. This analysis yields estimates of the transition point $p_{c}=0.49(1)$, the order-parameter critical exponent $\beta=0.16(2)$, and the correlation-length critical exponent $\nu=2.1(2)$ (see inset for collapsed curves). Error bars on these quantities are obtained as in Ref. [56], namely by examining where in the three-dimensional parameter space the $\chi^{2}$ cost function exceeds the minimum value by 30%.
@@
We now characterize the finite-size effects at the control transition by considering the small-size data in Fig. 4 (red data points), which corresponds to the system sizes for our simulations of the quantum limit of the model described in the next section. The order parameter for the small-system data shows a finite-size crossing near $p=0.65$, which naively would suggest that the order parameter remains finite at the transition and that the transition occurs at a value significantly larger than the one obtained from the large-system results. However, using the large-system results as a prior allows one to extract the critical properties from these data as well. In particular, we can analyze the small-system data assuming a scaling function of the form
$\overline{\braket{H_{ZZ}(t/L^{z}\!\to\!\infty)}}_{p=p_{c}}\!\!\sim\!\frac{1}{L^{\beta/\nu}}f\!\left[\!L^{1/\nu}(p-p_{c})\!+\!A_{C}L^{-\alpha}\!\right]$ (12)
@@
with the values of $\beta,\nu,$ and $p_{c}$ fixed to those extracted from the large-system data. $A_{C}$ and $\alpha$, which characterize the finite-size effects of the classical model, are then taken as fitting parameters for a separate $\chi^{2}$ analysis. Their optimal values $A_{C}\approx 9.46$ and $\alpha\approx 0.94$ collapse the small- and large-system data, as shown in the inset of Fig. 4. We note that the finite-size correction to the scaling variable, $A_{C}L^{-\alpha}$, takes values between $\sim 0.70$ ($L=16$) and $\sim 0.36$ for the small-system data (compared to $\lesssim 0.04$ for the large-system data), indicating strong finite-size effects at the system sizes accessible to simulations of the quantum limit of the model.
@@
Figure 5: Dynamics of the order parameter (classical model). Realization-averaged order parameter at $p=0.5$, near the control transition. Inset: Finite-size scaling collapse of the order-parameter dynamics yields a dynamical exponent $z=0.84(9)$. The red line shows a fit to a power-law decay $\sim 1/t^{0.104}$ which is close to the expected decay exponent $\beta/(\nu z)\approx 0.091$.
@@
The dynamical exponent $z$ of the control can be deduced from the time evolution of the order parameter near the transition. In Fig. 5, we show the dynamics of the order parameter for the large-system data at $p=0.5$. To determine the dynamical exponent, we collapse these data to a function of the form $L^{-\beta/\nu}f(t/L^{z})$, leading to a dynamical exponent $z=0.84(9)$ that produces the collapse shown in the inset. As a further self-consistency check, we note that combining the spacetime scaling $t\sim\xi^{z}$ with the scaling forms $\xi\sim|p-p_{c}|^{-\nu}$ and $\overline{\braket{H_{ZZ}(t/L^{z}\to\infty)}}_{p=p_{c}}\sim|p-p_{c}|^{\beta}$, we are led to the prediction that the order parameter at the transition should decay with time as a power law $\sim t^{\beta/(\nu z)}\approx t^{-0.091}$. The inset of Fig. 5 shows the collapsed curves on a log-log scale, with a red line indicating a best-fit power law obtained from the $L=540$ data. The best-fit power law $\sim t^{-0.104}$ is close to the value expected based on the above scaling argument.
@@
It is notable that a dynamical exponent $z<1$ is obtained from this analysis. An exponent $z=1$ is expected in transitions governed by a conformal field theory, as in the space-time random MIPT [56]. In adaptive circuits, absorbing-state transitions belonging to the directed-percolation and parity-conserving universality classes have $z\approx 1.581$ and $1.744$, respectively [57, 16, 17]. An exponent $z=2$, associated with diffusively spreading critical correlations governed by a random walk, has been observed in the Bernoulli map with both local [8, 9, 10] and global feedback [7, 10]. Infinite randomness [58] and quasiperiodic [59] fixed points with $z>1$ have also been identified in stabilizer circuits with static but spatially varying measurement probabilities. On the other hand, dynamical exponents $z<1$ have been found to occur in MIPTs with “power-law interacting” entangling gates [34] and at equilibrium phase transitions with sufficiently long-range power-law interactions [60] and are associated with superballistic spreading of correlations and sublinear power-law lightcones [61]. We attribute this feature of the transition to the sublattice magnetization measurements, which are themselves nonlocal, albeit spatially uniform.

### v2

@@ Abstract
Adaptive quantum circuits—where a quantum many-body state is controlled using measurements and conditional unitary operations—are a powerful paradigm for state preparation and quantum error correction tasks. They can support two types of nonequilibrium quantum phase transitions: measurement-induced transitions between volume- and area-law-entangled steady states and control-induced transitions where the system falls into an absorbing state or, more generally, an orbit visiting several absorbing states. Within this context, nonlocal conditional operations can alter the critical properties of the two transitions and the topology of the phase diagram. Here, we consider the scenario where the measurements are nonlocal, in order to engineer efficient control onto dynamical trajectories. Motivated by Rydberg-atom arrays, we consider a locally constrained model with global sublattice magnetization measurements and local correction operations to steer the system’s dynamics onto a many-body orbit with finite recurrence time. The model has a well-defined classical limit, which we leverage to aid our analysis of the control transition. As a function of the density of local correction operations, we find control and entanglement transitions with continuously varying critical exponents. For sufficiently high densities of local correction operations, we find that both transitions acquire a dynamical critical exponent $z<1$, reminiscent of criticality in long-range power-law interacting systems. At low correction densities, we find that the criticality reverts to a short-range nature with $z\gtrsim 1$. In the long-range regime, the control and entanglement transitions are indistinguishable to within the resolution of our finite-size numerics, while in the short-range regime we find evidence that the transitions become distinct. We conjecture that the effective long-range criticality mediated by collective measurements is essential in driving the two transitions together.
@@ I Introduction
With experimentally accessible noisy intermediate scale quantum (NISQ) devices, it is a timely and fundamental question to extend these ideas to monitored quantum systems. In a prototypical class of dynamical maps, Refs [3, 4] defined a probabilistic control protocol for the $\beta$-adic Rényi maps [6] which can also be used to control analogous quantum dynamics [7, 8, 9, 10]. The quantum control problem is enriched by measurements (necessary for the control protocol) which entail a backaction on the quantum state of the system, resulting in a phase transition in the late-time entanglement content of the quantum state. The entanglement transition can either precede or coincide with the (putatively) classical control transition depending on details of the entangling gates and control operation [11, 8, 10]. Thus, probabilistic control of quantum systems fits into the broader context of adaptive quantum dynamics [12, 13, 14, 15, 16, 17, 18, 19, 20, 11, 21, 22, 23, 24, 25, 26], where measurement and feedback can drive critical phenomena separately witnessed by local order parameters and entanglement measures [27, 28, 29, 30, 31, 32, 33]. This class of adaptive quantum dynamics holds promise for applications in quantum technologies ranging from robust state preparation to quantum error correction.
@@
The majority of monitored quantum circuits considered to date consist of short- to long-ranged unitary quantum gates and random local measurements. While measurements that project each local qubit fully destroy the entanglement of the many-body wavefunction, it is possible to perform a collective measurement of a global property of the ensemble of spins (e.g., their total magnetization) which has an extensive number of possible measurement outcomes and does not fully disentangle the state. Long-range gates have a strong effect on the universality class of the measurement induced phase transition (MIPT) [34, 35, 36], but it remains unclear if such a transition can persist in the presence of long-range or collective measurements and if so, what is its universal nature. Additionally, these collective measurements allows for a wider range of adaptive dynamics where we can feedback the results of these measurements to access a wider range fixed-point structures which could be used to overcome strongly entangled dynamics.
@@
In this paper, we utilize adaptive quantum circuits as a framework to explore control and entanglement transitions in a setting that requires collective measurements; we explore this in a model of probabilistic quantum dynamics inspired by Rydberg-atom quantum simulators [37, 38]. The model consists of a chaotic circuit driven by kinetically constrained Rydberg-blockade dynamics [39, 40, 41, 42] that is stochastically interleaved with collective measurements and local control protocol (see Fig. 1). The chaotic circuit (defined by $p=0$) possesses an unstable Hilbert-space orbit where the system cycles through a finite number of product states [43, 44, 39]. The control protocol attempts to convert this unstable orbit into an attractor, i.e., a steady state that is ultimately reached regardless of the initial state, using collective measurements and local correction operations. While our control protocol is applied stochastically, the dynamics within a protocol are necessarily adaptive. We demonstrate the existence of control and entanglement transitions in this model with the aid of a classical limit in which the model maps onto a classical probabilistic cellular automaton (CA); this limit manifests a many-body generalization of classical probabilistic control. With a variety of metrics, we estimate the locations and critical exponents of both transitions and find evidence that, in a large portion of the phase diagram, the transitions coincide and belong to the same universality class (although we cannot rule out the possibility that the transitions occur quite close to one another and with similar critical exponents). Intriguingly, for a wide range of parameters, both transitions possess a dynamical critical exponent $z<1$, similar to MIPTs in monitored quantum dynamics with long-range unitary gates [34].
@@
The remainder of the paper is organized as follows. In Sec. II.1, we define the probabilistic control model and its classical limit. In Sec. III, we study the control transition in the quantum model for a particular cut through the phase diagram and show where its critical properties match the transition in the classical model. In Sec. IV, we study the entanglement transition along the same cut through the lens of entanglement entropy, mutual information, and purification dynamics. We consider the phase diagram in Sec. V, and conclude in Sec. VI with a discussion and outlook. Appendix A discusses how to implement the collective measurements underlying our control protocol in a qubit-only setup. Appendices B and C discuss generalizations of our control protocol and the impact of finite-size effects, respectively.
@@ II.2 Control Protocol
Note that this measurement does not collapse |$\ket{\psi}$ to a single CB state. In fact, the subspace $\mathcal{H}_{m_{\rm even(odd)}}$ contains exponentially many CB states and therefore can support volume-law entanglement. For this reason, we refer to this type of measurement, which reveals global information about the state without fully collapsing it to a product state, as a collective measurement. Such a measurement can be implemented experimentally in a qubit-only setup by computing the magnetization onto a register of $O(\log_{2}L)$ ancilla qubits [52, 53, 54, 55], or in a circuit/cavity QED setup by dispersively coupling the qubits in each sublattice to a separate cavity mode and measuring the frequency shift of either mode. We illustrate one explicit way to perform this measurement with ancillary qubits in Appendix A. Such collective measurements have been achieved (mid-circuit) in arrays of tweezer-trapped Rydberg atoms coupled to cavities [56, 57].
@@
is applied on site $k=j$ or $j+2$. $U^{c}_{k}$ flips qubit $k$ provided the flip does not violate the Fibonacci constraint described below Eq. (1). The table at the bottom of Fig. 3(b) shows whether qubit $j$ or $j+2$ is flipped, depending on the target state and the outcome of the $Z_{j}$ measurement. In Appendix B, we discuss how this control protocol can be generalized to other Hilbert space orbits with finite recurrence times, of the sort discussed in Ref. [39].
Both the collective measurements and the resulting feedback are crucial for directing the system to one of the three vacuum states. The collective measurements find the “closest” orbit, and the feedback adapts based on the measurement result to direct the system closer to the target state.
The strength of the control is quantified by two parameters: the control rate $p$ that sets the probability with which control is applied at a given time step, and the probability $q$ that sets the fraction of sites to which the local feedback operation is applied. The addition of the “spatial” probability $q$ is particularly important in that it represents a step beyond the probabilistic control literature [3, 2, 4, 5] that has so far focused on single-body chaotic dynamics  [^1] In both cases, $p$ and $q$ are classical probabilities implementable, for instance, with a pseudo-random number generator on a classical computer.. In generalizing to the many-body setting, we allow for the possibility that only a finite fraction of sites are subjected to control (beyond the initial step in which some global information about the system’s state is measured).
@@ II.3 Classical Limit
Both the chaotic circuit and the control protocol have a numerically tractable classical limit that is useful in our analysis of the control transition. In this limit, the dynamics of an initial CB state does not lead to superposition and entanglement. Rather, the system is always in a CB state |$\ket{b}$, corresponding to a bitstring $b$. At each time step, the CB state |$\ket{b}\mapsto\ket{b^{\prime}}$ up to a global phase. Thus, to keep track of the dynamics of an initial CB state and the evolution of observables that are diagonal in the CB [like the control order parameter $H_{\rm ZZ}$, Eq. (3)], it is sufficient to follow the dynamics of the bitstring $\dots\mapsto b\mapsto b^{\prime}\mapsto\dots$. The dynamics in the classical limit is therefore equivalent to a CA. This is clearly visible in Fig. 2(b,d,f) where the system always has expectation value $\braket{Z_{j}}=\pm 1$ on each site and time, indicating that the system remains in a CB state throughout the dynamics.
@@ III Control Transition
We now consider the existence of a control transition in this model, before turning our focus in Sec. IV to the interplay of control and entanglement properties. Previous work on $\beta$-adic Rényi circuits [7, 8, 9, 10] suggests that properties of the control transition are primarily governed by classical physics. For this reason, we first consider the automaton limit of our model [$\theta=\pi/2$ in Eq. (4)], whose classical simulability gives access to larger system sizes and better estimates of critical properties that can serve as a useful guide when analyzing the quantum limit. In our numerical simulations, we set $q=0.2$ and measure the control order parameter $H_{ZZ}$ after evolving the system for $L$ time steps from an initial CB state. (We defer discussion of the $q$-dependence of our results to Sec. V.) Results are averaged over $10^{4}$ samples comprising randomly chosen initial CB states, circuit realizations, and quantum measurement outcomes (sampled according to their Born probabilities). Each sample yields an expectation value of the control order parameter that we denote $\braket{H_{ZZ}(t)}=\langle\psi(t)|H_{ZZ}|\psi(t)\rangle$. In the classical limit, $\braket{H_{ZZ}(t)}$ takes discrete values between 00 and $1$ in steps of $1/L$ since the system is always in a CB state; in the quantum limit, the system can be in a superposition of CB states and $\braket{H_{ZZ}(t)}$ becomes a continuous variable. The sample average $\overline{\braket{H_{ZZ}(t)}}$, along with error bars denoting the standard error of the mean, is the object of our analysis below.
@@
Before proceeding, we offer a remark about finite-size effects. The control protocol defined in Sec. II.2 targets the vacuum orbit Eq. (2), which is invariant under the chaotic circuit defined in Sec. II.1. If the control protocol successfully drives the system onto the orbit, it can never escape. In a finite-size system, even at low control rate there is a finite probability that the system becomes trapped on the orbit. This probability is exponentially small in the system size $L$, but this means that such rare events can occur on timescales $\sim e^{O(L)}$. At large system sizes like those accessible in the classical limit, this is not an issue. However, at small system sizes like those accessible in the quantum limit, this finite-size effect manifests as a late-time decay of the control order parameter, even in the chaotic phase where one would expect it to saturate to a time-independent value. We discuss the decay further in Appendix C and provide a mechanism to remove it if desired. Nevertheless, we will see that it is still possible to perform a consistent finite-size scaling analysis in the presence of this decay  [^2] We note that this effect is also present in studies of absorbing state transitions in monitored quantum dynamics [15, 17, 16], but was not present in studies of probabilistic control in $\beta$-adic Rényi circuits [7, 8, 9, 10], where a random perturbation was used to remove the orbits while preserving the control transition..
@@ III.1 Classical Control Transition
Figure 4: Order parameter (classical model). Shown at time $t=L^{0.86}$ as a function of control rate $p$ in the classical limit of the model for both large (blue) and small (red) system sizes. Here and in all subsequent figures, error bars indicating the standard error of the mean over $10^{4}$ samples are present, but are smaller than the plot markers. Inset: Finite-size scaling collapse of the large-system data indicates $p_{c}=0.49(1)$, $\beta=0.17(2)$, and $\nu=2.3(2)$. Small-system data are collapsed assuming a scaling function of the form $L^{-\beta/\nu}f[L^{1/\nu}(p-p_{c})+AL^{-\alpha}]$ with $p_{c}$, $\beta$, and $\nu$ fixed by the large-system data collapse. This yields $A\approx 9.61$ and $\alpha\approx 1.08$, which quantify the finite-size effects present at system sizes accessible to the quantum numerics.
@@
In Fig. 4, we plot the sample-averaged order parameter $\overline{\braket{H_{ZZ}(t=L^{0.86})}}$  [^3] Note that, since $z$ is not generically an integer, evaluating observables at time $t=c\,L^{z}$ for some $O(1)$ constant $c$ requires obtaining $\overline{S_{\rm anc}}$ at noninteger times. In this work we use a linear interpolation to estimate the values of observables at $t=c\,L^{z}$ and to determine the associated error bars from those of the neighboring data points. for a range of $p$ and two families of system sizes: $L=300,\dots,540$ (blue points) and $L=16,\dots,32$ (red points). The exponent $0.86$, which yields the best scaling collapse under the protocol described in the next paragraph, is an estimate of the dynamical critical exponent $z$. From these data, we extract an estimate of the critical point $p_{c}$. The dynamics near $p_{c}$, shown in Fig. 5, are then analyzed to yield a refined estimate $z=0.8(1)$ which is within error bars of the value used to obtain Fig. 4. The large-system-size data will be used to extract the location and critical exponents of the control transition. The small-system-size data represent system sizes accessible to our fully quantum simulations that will be reported in the next subsection. We first describe the transition from the perspective of the large-system data before returning to the small-system data to analyze the finite-size effects that will be important when considering the quantum data.
@@
To extract the critical point $p_{c}$, we perform finite-size scaling collapse assuming a scaling ansatz $\overline{\braket{H_{ZZ}(t=L^{z})}}=L^{-\beta/\nu}f[L^{1/\nu}(p-p_{c})]$. This corresponds to a second-order phase transition where the order parameter vanishes near the transition as $\overline{\braket{H_{ZZ}(t=L^{z})}}_{p\sim p_{c}}\sim|p-p_{c}|^{\beta}$ (in the thermodynamic limit) and the correlation length diverges as $\xi\sim|p-p_{c}|^{-\nu}$. The scaling collapse (like all others reported in this paper) is performed using a standard $\chi^{2}$ analysis [61] that uses a numerical least-squares opimization to collapse the curves. This analysis yields estimates of the transition point $p_{c}=0.49(1)$, the order-parameter critical exponent $\beta=0.17(2)$, and the correlation-length critical exponent $\nu=2.3(2)$ (see inset for collapsed curves). Error bars on these quantities are obtained as in Ref. [62], namely by examining where in the three-dimensional parameter space the $\chi^{2}$ cost function exceeds the minimum value by 30%.
@@
We now characterize the finite-size effects at the control transition by considering the small-size data in Fig. 4 (red data points), which corresponds to the system sizes for our simulations of the quantum limit of the model described in the next section. The order parameter for the small-system data shows a finite-size crossing near $p=0.7$, which naively would suggest that the order parameter remains finite at the transition and that the transition occurs at a value significantly larger than the one obtained from the large-system results. However, using the large-system results as a prior allows one to extract the critical properties from these data as well. In particular, we can analyze the small-system data assuming a scaling function of the form
$\overline{\braket{H_{ZZ}(t=L^{z})}}_{p>p_{c}}\!\!\sim\!\frac{1}{L^{\beta/\nu}}f\!\left[\!L^{1/\nu}(p-p_{c})\!+\!A_{C}L^{-\alpha}\!\right]$ (12)
@@
with the values of $\beta,\nu,$ and $p_{c}$ fixed to those extracted from the large-system data. $A_{C}$ and $\alpha$, which characterize the finite-size effects of the classical model, are then taken as fitting parameters for a separate $\chi^{2}$ analysis. Their optimal values $A_{C}\approx 9.61$ and $\alpha\approx 1.08$ collapse the small- and large-system data, as shown in the inset of Fig. 4. We note that the finite-size correction to the scaling variable, $A_{C}L^{-\alpha}$, takes values between $\sim 0.48$ ($L=16$) and $\sim 0.23$ ($L=32$) for the small-system data (compared to $\lesssim 0.02$ for the large-system data), indicating strong finite-size effects at the system sizes accessible to simulations of the quantum limit of the model.
@@
Figure 5: Dynamics of the order parameter (classical model). Realization-averaged order parameter at $p=0.5$, near the control transition. Inset: Finite-size scaling collapse of the order-parameter dynamics yields a dynamical exponent $z=0.8(1)$. The red line shows a fit to a power-law decay $\sim 1/t^{0.104}$ which is close to the expected decay exponent $\beta/(\nu z)\approx 0.093$.
@@
The dynamical exponent $z$ of the control can be deduced from the time evolution of the order parameter near the transition. In Fig. 5, we show the dynamics of the order parameter for the large-system data at $p=0.5$. To determine the dynamical exponent, we collapse these data to a function of the form $L^{-\beta/\nu}f(t/L^{z})$ with $\beta/\nu$ fixed by the analysis in the previous paragraph. This leads to an estimated dynamical exponent $z=0.8(1)$ that produces the collapse shown in the inset. As a further self-consistency check, we note that combining the spacetime scaling $t\sim\xi^{z}$ with the scaling forms $\xi\sim|p-p_{c}|^{-\nu}$ and $\overline{\braket{H_{ZZ}(t=L^{z})}}_{p\sim p_{c}}\sim|p-p_{c}|^{\beta}$, we are led to the prediction that the order parameter should decay with time as a power law $\sim t^{\beta/(\nu z)}\approx t^{-0.093}$ near the transition. The inset of Fig. 5 shows the collapsed curves on a log-log scale, with a red line indicating a best-fit power law obtained from the $L=540$ data. The best-fit power law $\sim t^{-0.104}$ is close to the value expected based on the above scaling argument.
@@
It is notable that a dynamical exponent $z<1$ is obtained from this analysis. An exponent $z=1$ is expected in transitions governed by a conformal field theory, as in the space-time random MIPT [62]. In adaptive circuits, absorbing-state transitions belonging to the directed-percolation and parity-conserving universality classes have $z\approx 1.581$ and $1.744$, respectively [63, 16, 17]. (Further, recent work shows a family that interpolates between these classes [64].) An exponent $z=2$, associated with diffusively spreading critical correlations governed by a random walk, has been observed in the Bernoulli map with both local [8, 9, 10] and global feedback [7, 10]. Infinite randomness [65] and quasiperiodic [66] fixed points with $z>1$ have also been identified in stabilizer circuits with static but spatially varying measurement probabilities. On the other hand, dynamical exponents $z<1$ have been found to occur in MIPTs with “power-law interacting” entangling gates [34] and at equilibrium phase transitions with sufficiently long-range power-law interactions [67] and are associated with superballistic spreading of correlations and sublinear power-law lightcones [68]. We attribute this feature of the transition to the sublattice magnetization measurements, which are themselves nonlocal, albeit spatially uniform.